# A Survey of Graph Retrieval-Augmented Generation for Customized Large Language Models

### v1

@@ Abstract
Retrieval-Augmented Generation, Knowledge Graphs, Large Language Models, GraphRAG
@@ I Introduction
Large language models (LLMs), like the GPT series [1], have surprised the world with their remarkable ability across a wide range of tasks, achieving breakthroughs in text comprehension [2], question answering [3], and content generation [4, 5, 6]. Despite their effectiveness, LLMs are always criticized for their limited ability to handle knowledge-intensive tasks, especially when faced with questions requiring domain expertise [7]. Specifically, the application of LLMs to specialized domains remains challenging for three fundamental reasons, including ❶ Knowledge limitations: LLMs’ pre-trained knowledge is broad but shallow in specialized fields. Their training data primarily consists of general-domain content, leading to insufficient depth in professional domains and potential inconsistencies with current domain-specific standards and practices. ❷ Reasoning complexity: Specialized domains require precise, multi-step reasoning with domain-specific rules and constraints. LLMs often struggle to maintain logical consistency and professional accuracy throughout extended reasoning chains, particularly when dealing with technical constraints or domain-specific protocols. ❸ Context sensitivity: Professional fields often involve context-dependent interpretations where the same terms or concepts may have different meanings or implications based on specific circumstances. LLMs frequently fail to capture these nuanced contextual variations, leading to potential misinterpretations or inappropriate generalizations.
@@
Retrieval-Augmented generation (RAG) offers a promising solution to customize LLMs for specific domains [12]. Rather than retraining LLMs to incorporate updates, RAG enhances these models by leveraging external knowledge from text corpora without modifying their architecture or parameters. This approach enables LLMs to generate responses by leveraging not only their pre-trained knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers. The naive RAG systems operate through three key steps: knowledge preparation, retrieval, and integration. During knowledge preparation, external sources such as documents, databases, or webpages are divided into manageable textual chunks and converted into vector representations for efficient indexing. In the retrieval stage, when a user submits a query, the system searches for relevant chunks using keyword matching or vector similarity measures. The integration stage then combines these retrieved chunks with the original query to create an informed prompt for the LLM’s response. Recently, some advanced RAG systems have evolved beyond simple text chunk retrieval to offer more sophisticated knowledge augmentation approaches. These include hierarchical RAG that preserves document structure through multi-level retrieval [13, 14], re-ranking systems that implement two-stage retrieval for higher recall and precision [15, 16], self-querying RAG that automatically decomposes complex queries [17], and adaptive RAG that dynamically adjusts retrieval strategies based on query types [18, 19]. These advanced RAG systems aim to address the limitations of naive RAG approaches by improving context awareness, retrieval accuracy, and handling complex queries more effectively.
@@
The last challenge relates to system efficiency and scalability. The entire RAG pipeline - from initial corpus preprocessing and indexing to real-time retrieval and generation - faces significant efficiency bottlenecks [25, 26]. The external knowledge base contains a lot of domain-irrelevant information, while domain-specific terminologies are always sparsely distributed over these documents. RAG systems can be computationally expensive and time-consuming [25], especially when dealing with large-scale knowledge sources, as the model needs to search through vast amounts of unstructured text to find relevant information. Moreover, real-time retrieval and cross-document reasoning can introduce considerable latency, negatively impacting user experience. The scalability of RAG is further constrained by declining retrieval quality and accuracy as the size of the knowledge base grows [26], thereby limiting its practical deployment in extensive and dynamic professional environments.
@@
To address these limitations, graph retrieval-augmented generation (GraphRAG) has recently emerged as a new paradigm to customize LLMs with well-organized background knowledge and improved contextual reasoning [25, 27, 28, 29]. Based on the utilization of graph structures, existing GraphRAG models can be categorized into three main categories: ❶ Knowledge-based GraphRAG, which uses graphs as knowledge carriers. ❷ Index-based GraphRAG uses graphs as index tools to retrieve relevant raw texts from the corpus, and ❸ Hybrid GraphRAG which combines the strengths of both knowledge-based and index-based frameworks, providing more advanced solutions for complex reasoning tasks. Knowledge-based and Index-based GraphRAG represent two distinct paradigms for enhancing LLMs with structured knowledge. Knowledge-based GraphRAG focuses on transforming unstructured textual documents into explicit and structured KGs, where nodes represent domain concepts and edges capture semantic relationships between them, enabling better representation of hierarchical relationships and complex knowledge dependencies. In contrast, Index-based GraphRAG maintains the original textual form while utilizing graph structures primarily as an indexing mechanism to organize and retrieve relevant text chunks efficiently. By incorporating graph structures into text indexing, Index-based GraphRAG methods establish semantic connections between text chunks for efficient look-up operations and retrieval. While Knowledge-based GraphRAG emphasizes the explicit modeling of domain knowledge and semantic relationships through graph transformation, Index-based GraphRAG prioritizes efficient information retrieval and global navigation through the graph-based organization of the raw text. This fundamental difference in approach reflects their distinct purposes: Knowledge-based GraphRAG aims to create a structured knowledge representation for a better understanding of complex relationships with graph-based reasoning capability, whereas Index-based GraphRAG focuses on optimizing the retrieval and accessibility of relevant textual information through graph-based indexing strategies.
@@
Figure 2: A comprehensive overview of traditional RAG and two typical GraphRAG workflows. Non-graph RAG organizes the corpus into chunks, ranks them by similarity, and retrieves the most relevant text for generating responses. Knowledge-based GraphRAG extracts detailed knowledge graphs from the corpus using entity recognition and relation extraction, offering fine-grained, domain-specific information. Index-based GraphRAG summarizes the corpus into high-level topic nodes, which are linked to form an index graph, while fact linking maps topics to text. This two-layer structure combines efficient topic retrieval with detailed text knowledge, offering scalability and performance compared to Knowledge-based GraphRAG.
@@ II Overview
This survey provides a comprehensive analysis of GraphRAG, detailing its taxonomy, mechanisms, challenges, and future research directions, which is organized into seven main sections that progress from foundational concepts to practical implementations. Specifically, we begin in Section 2 II by establishing the foundational framework, tracing GraphRAG’s evolution from traditional RAG systems, examining RAG’s limitations in handling structured knowledge, and introducing GraphRAG’s core concepts and advantages in complex reasoning tasks. The following three sections systematically explore the key components of GraphRAG systems: two primary paradigms for structured knowledge organization, including knowledge carrier graphs and index graphs (Section 3 IV), retrieval techniques for extracting query-relevant factual information from structured knowledge bases (Section 4 V), and knowledge integration methods for effectively incorporating retrieved knowledge into LLMs (Section 5 VI). Moving toward practical applications, Section 6 VIII addresses implementation aspects by providing detailed guidelines, reviewing open-source projects, and presenting domain-specific case studies supported by comprehensive datasets and evaluation benchmarks. Finally, Section 7 VII concludes by identifying future research directions and discussing potential challenges in knowledge quality, retrieval efficiency, system generalization, and security, alongside practical guidance for building domain-specific GraphRAG systems. Throughout the survey, we balance fundamental concepts, the current state of the art, and practical implementations, making it valuable for researchers advancing the field and practitioners developing GraphRAG applications in real-world scenarios.
@@
Our survey advances beyond the existing survey [28, 29, 30] through a more systematic and comprehensive approach to GraphRAG systems. While the previous survey offered a basic workflow description covering Graph-based Indexing, Graph-guided Retrieval, and Graph-enhanced Generation, we introduce a more sophisticated and comprehensive taxonomy that clearly categorizes GraphRAG approaches into three distinct categories (Knowledge-based, Index-based, and Hybrid GraphRAG), providing a more nuanced understanding of the field. Our survey features a more systematic six-section structure that progresses logically from theoretical foundations to practical implementations, offering a more detailed exploration of each component, including knowledge organization paradigms, retrieval techniques, and integration methods. Unlike the previous survey, we provide extensive practical guidance through detailed review of open-source projects, and domain-specific case studies supported by comprehensive datasets and evaluation benchmarks. We also offer a more thorough analysis of challenges and solutions across multiple dimensions, including knowledge quality, retrieval efficiency, system generalization, and security concerns. Finally, while the existing survey broadly discussed potential applications, we provide actionable insights supported by empirical evidence and implementation examples, making our survey a more valuable resource for practitioners deploying GraphRAG systems in production environments.
@@
Indexing, leaf-datasets, text width=6em [GNN-ret [31], PG-RAG [32], KGP [33], GraphCoder [34], AVIS[35], SayPlan [36] HotpotQA [37], GraphRetriever [38], modelnode-datasets, text width=50.7em] ] [ Graphs as
@@
Construction from Corpus, leaf-datasets, text width=11em [StructRAG [14], GraphRAG [25], LightRAG [26], QUEST [39], GraphReader [40],
Structure-guided Prompts [41], modelnode-datasets, text width=38em] ] [ GraphRAG with Existing
@@
SubGraphRAG [46], modelnode-datasets, text width=38em] ] ] [ Hybrid
GraphRAG, leaf-datasets, text width=6em [GoR [47], MedGraphRAG [48], CodexGraph [49], modelnode-datasets, text width=50.7em] ] ] [ Knowledge
@@
CancerKG [51], G-Retriever [52], modelnode-metrics, text width=38em] ] [ Logical-based Retriever, leaf-metrics, text width=11em [RoG [42], RD-P [53], RuleRAG [54], KGL [55], RiTeK [56], modelnode-metrics, text width=38em] ] [ GNN-based Retriever, leaf-metrics, text width=11em [GNN-Ret [31], SURGE [57], GNN-RAG [58], modelnode-metrics, text width=38em] ] [ LLM-based Retriever, leaf-metrics, text width=11em [ToG [22], Graph RAG [25], LightRAG [26], KGP [33], MEG [59], TQA-KG [60], modelnode-metrics, text width=38em] ] [ RL-based Retriever, leaf-metrics, text width=11em [KnowGPT [7], Spider [61], modelnode-metrics, text width=38em] ] ] [ Retrieval
Strategy, leaf-metrics, text width=6em [ Multi-round Retrieval, leaf-metrics, text width=11em [GoR [47], DialogGSR [62], Graph-CoT [63], modelnode-metrics, text width=38em] ] [ Post-retrieval, leaf-metrics, text width=11em [KGR [43], CoK [64], modelnode-metrics, text width=38em] ] [ Hybrid Retrieval, leaf-metrics, text width=11em [StructRAG [14], ToG 2.0 [23], modelnode-metrics, text width=38em] ] ] ] [ Knowledge
@@
Knowledge, leaf-methods, text width=11em [RHO [70], GNP [71], InstructGLM [72], LLAGA [73], modelnode-methods, text width=38em] ] ] [ In-context
@@
Graph Refinement, leaf-methods, text width=11em [KELP [44], FMEA-KG [78], EtD [79], CogMG [80], modelnode-methods, text width=38em] ] ] ] ]
@@
III What is GraphRAG
III-A Traditional RAG Pipeline
A RAG framework begins by retrieving relevant information from pre-constructed external knowledge bases based on the query. This information is then used to prompt LLMs, guiding them in constructing credible reasoning chains. As a result, RAG enables LLMs to generate more substantiated and accurate content, effectively minimizing hallucinations and inconsistencies. The traditional RAG pipeline typically comprises three core components: knowledge organization, knowledge retrieval, and knowledge integration.
III-A1 Knowledge organization
In summary, knowledge organization is foundational to the retrieval process. By carefully constructing the knowledge resources, RAG systems can ensure the fidelity and reliability of retrieved content.
III-A2 Knowledge retrieval
Current RAG works usually involve retrieval methods such as k-nearest neighbor retrieval (KNN), term frequency-inverse document frequency (TF-IDF), and best matching 25 (BM25) to retrieve the relevant content. RETRO [81] employs KNN to extract approximate relevant neighbors from the conducted key-value database by calculating the L2 distance. RETROprompt [90] extends this approach to construct a few-shot knowledge store, tailoring it specifically for prompt learning scenarios.
Retrieval model training is another crucial aspect, where recent methodologies have explored the use of self-supervised techniques. For instance, REPLUG [94] demonstrates that labels generated by a frozen language model can be leveraged to directly supervise the training of retrieval models. This approach enhances retrieval quality without requiring additional manually annotated datasets. ATLAS [82] fine-tunes both the retriever and the LLMs in tandem. Furthermore, frameworks like FLARE [83] dynamically balance retrieval with generation needs through an active RAG framework that intelligently determines retrieval timings based on the generation progression.
Moreover, the use of external APIs helps to broaden the scope of retrieval capabilities. By integrating external knowledge sources like Wikipedia and Google Search APIs, RAG systems can access expansive real-time databases, enriching their generated content. For example, Toolformer [95] taps into Wikipedia to access extensive knowledge bases, while the work by Lazaridou et al. [96] utilizes Google Search to ensure the most current and comprehensive information is incorporated, granting RAG systems a dynamic edge in content generation.
Overall, the knowledge retrieval process combines innovative methodologies with diverse data sourcing strategies, underpinning the generation of informed and contextually relevant outputs.

### v2

@@ Abstract
Retrieval-Augmented Generation, Knowledge Graphs, Graph-based Retrieval-Augmented Generation
@@ I Introduction
Large language models (LLMs), like the GPT series [1], have surprised the world with their remarkable ability across a wide range of tasks, achieving breakthroughs in text comprehension [2], question answering [3], and content generation [4]. Despite their effectiveness, LLMs are always criticized for their limited ability to handle knowledge-intensive tasks, especially when faced with questions requiring domain expertise [5]. Specifically, applying LLMs into specialized domains remains challenging for three fundamental reasons, including ❶ Knowledge limitations: LLMs’ pre-trained knowledge is broad but shallow in specialized fields. Their training data primarily consists of general-domain content, leading to insufficient depth in professional domains and potential inconsistencies with current domain-specific standards and practices. ❷ Reasoning complexity: Specialized domains require precise, multi-step reasoning with domain-specific rules and constraints. LLMs often struggle to maintain logical consistency and professional accuracy throughout extended reasoning chains, particularly when dealing with technical constraints or domain-specific protocols. ❸ Context sensitivity: Professional fields often involve context-dependent interpretations where the same terms or concepts may have different meanings or implications based on specific circumstances. LLMs frequently fail to capture these nuanced contextual variations, leading to potential misinterpretations or inappropriate generalizations.
@@
Retrieval-augmented generation (RAG) offers a promising solution to customize LLMs for specific domains [10, 11]. Rather than retraining LLMs to incorporate updates, RAG enhances these models by leveraging external knowledge from text corpora without modifying their architecture or parameters. This approach enables LLMs to generate responses by leveraging not only their pre-trained knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers. The naive RAG systems operate through three key steps: knowledge preparation, retrieval, and integration. During knowledge preparation, the external textual corpus is divided into manageable textual chunks and converted into vector representations for efficient indexing. In the retrieval stage, when a user submits a query, the system searches for relevant chunks using keyword matching or vector similarity measures. The integration stage then combines these retrieved chunks with the original query to generate more informed responses. Recently, some advanced RAG systems have evolved beyond simple text chunk retrieval to offer more sophisticated knowledge augmentation approaches. These include hierarchical RAG that preserves document structure through multi-level retrieval [12, 13], re-ranking systems that implement two-stage retrieval for higher recall and precision [14, 15], self-querying RAG that automatically decomposes complex queries [16], and adaptive RAG that dynamically adjusts retrieval strategies based on query types [17, 18]. These strategies advance naive RAG systems by improving context awareness, retrieval accuracy, and handling complex queries.
@@
The last challenge relates to system efficiency and scalability. The entire RAG pipeline - from initial corpus preprocessing and indexing to real-time retrieval and generation - faces significant efficiency bottlenecks [24, 25]. The external knowledge base contains a lot of domain-irrelevant information, while domain-specific terminologies are always sparsely distributed over these documents. RAG systems can be computationally expensive and time-consuming [24], especially when dealing with large-scale knowledge sources, as the model needs to search through vast amounts of unstructured text to find relevant information. Moreover, real-time retrieval and cross-document reasoning can introduce considerable latency, negatively impacting user experience. The scalability of RAG is further constrained by declining retrieval quality and accuracy as the size of the knowledge base grows [25], thereby limiting its practical deployment in extensive and dynamic environments.
@@
To address these limitations, graph retrieval-augmented generation (GraphRAG) has recently emerged as a new paradigm to customize LLMs with well-organized background knowledge and improved contextual reasoning [24, 26, 27, 28, 29]. Based on the utilization of graph structures, existing GraphRAG models can be categorized into three main categories: ❶ Knowledge-based GraphRAG, which uses graphs as knowledge carriers. ❷ Index-based GraphRAG uses graphs as index tools to retrieve relevant raw texts from the corpus, and ❸ Hybrid GraphRAG which combines the strengths of both knowledge-based and index-based frameworks, providing more advanced solutions for complex reasoning tasks. Knowledge-based and Index-based GraphRAG represent two distinct paradigms for enhancing LLMs with structured knowledge. Knowledge-based GraphRAG focuses on transforming unstructured textual documents into explicit and structured KGs, where nodes represent domain concepts and edges capture semantic relationships between them, enabling better representation of hierarchical relationships and complex knowledge dependencies. In contrast, Index-based GraphRAG maintains the original textual form while utilizing graph structures primarily as an indexing mechanism to organize and retrieve relevant text chunks efficiently. By incorporating graph structures into text indexing, Index-based GraphRAG methods establish semantic connections between text chunks for efficient look-up operations and retrieval. While Knowledge-based GraphRAG emphasizes the explicit modeling of domain knowledge and semantic relationships through graph transformation, Index-based GraphRAG prioritizes efficient information retrieval and global navigation through the graph-based organization of the raw text. This fundamental difference in approach reflects their distinct purposes: Knowledge-based GraphRAG aims to create a structured knowledge representation for a better understanding of complex relationships with graph-based reasoning capability, whereas Index-based GraphRAG focuses on optimizing the retrieval and accessibility of relevant textual information through graph-based indexing strategies.
@@
Figure 2: A comprehensive overview of traditional RAG and two typical GraphRAG workflows. Non-graph RAG organizes the corpus into chunks, ranks them by similarity, and retrieves the most relevant text for generating responses. Knowledge-based GraphRAG extracts detailed knowledge graphs from the corpus using entity recognition and relation extraction, offering fine-grained, domain-specific information. Index-based GraphRAG summarizes the corpus into high-level topic nodes, which are linked to form an index graph while fact-linking maps topics to text. This two-layer structure combines efficient topic retrieval with detailed text knowledge, offering scalability and performance compared to Knowledge-based GraphRAG.
@@ II Overview
This survey provides a comprehensive analysis of GraphRAG, detailing its taxonomy, mechanisms, challenges, and future research directions, which is organized into seven main sections that progress from foundational concepts to practical implementations. Specifically, we begin in Section III by establishing the foundational framework, tracing GraphRAG’s evolution from traditional RAG systems, examining RAG’s limitations in handling structured knowledge, and introducing GraphRAG’s core concepts and advantages in complex reasoning tasks. The following three sections systematically explore the key components of GraphRAG systems: two primary paradigms for structured knowledge organization, including knowledge carrier graphs and index graphs (Section IV), retrieval techniques for extracting query-relevant factual information from structured knowledge bases (Section V), and knowledge integration methods for effectively incorporating retrieved knowledge into LLMs (Section VI). Moving toward practical applications, Section -A (see Appendix) explores the real-world implementations of GraphRAG systems by reviewing open-source projects and benchmark datasets across different domains. Finally, Sections -B and -C (see Appendix) conclude by identifying future research directions and discussing potential challenges in knowledge quality, retrieval efficiency, system generalization, and security. Throughout the survey, we balance fundamental concepts, the current state-of-the-art methods, and practical implementations, making it valuable for both researchers advancing the field and practitioners developing GraphRAG applications in real-world scenarios.
@@
Our survey advances beyond the existing survey [27, 29, 30] through a more systematic and comprehensive review of existing GraphRAG systems. While the previous survey offered a basic workflow description covering Graph-based Indexing, Graph-guided Retrieval, and Graph-enhanced Generation, we introduce a more sophisticated and comprehensive taxonomy that clearly categorizes GraphRAG approaches into three distinct categories (Knowledge-based, Index-based, and Hybrid GraphRAG), providing a more nuanced understanding of the field. Our survey features a more systematic six-section structure that progresses logically from theoretical foundations to practical implementations, offering a more detailed exploration of each component, including knowledge organization paradigms, retrieval techniques, and integration methods. Unlike the previous survey, we provide extensive practical guidance through a detailed review of open-source projects, and domain-specific case studies supported by comprehensive datasets across different domains. We also offer a more thorough analysis of challenges and solutions across multiple dimensions, including knowledge quality, retrieval efficiency, system generalization, and security concerns. Finally, while the existing survey broadly discussed potential applications, we provide actionable insights supported by comprehensive analysis and implementation examples, making our survey a more valuable resource for practitioners to deploy GraphRAG systems in production environments.
III What is GraphRAG
III-A Traditional RAG Pipeline
A RAG framework begins by retrieving relevant information from pre-constructed external knowledge bases based on the query. This information is then used to prompt LLMs, guiding them in constructing credible reasoning chains. As a result, RAG enables LLMs to generate more substantiated and accurate content, effectively minimizing hallucinations and inconsistencies. The traditional RAG pipeline typically comprises three core components: knowledge organization, knowledge retrieval, and knowledge integration. We provide a detailed analysis of traditional RAG systems in Appendix D.
III-A1 Knowledge organization
In summary, knowledge organization is foundational to the retrieval process. By carefully constructing the knowledge resources, RAG systems can improve the retrieval efficiency and ensure the fidelity and reliability of retrieved content.
III-A2 Knowledge retrieval
Current RAG systems usually involve retrieval methods such as k-nearest neighbor retrieval (KNN), term frequency-inverse document frequency (TF-IDF), and best matching 25 (BM25) to retrieve the relevant content. RETRO [31] employs KNN to extract approximate relevant neighbors from the conducted key-value database by calculating the L2 distance. RETROprompt [40] extends this approach with a few-shot knowledge store, tailoring it for more advanced prompts.
Overall, knowledge retrieval combines innovative methodologies with diverse data-sourcing strategies, underpinning the generation of informed and contextually relevant outputs.
@@ III-A2 Knowledge retrieval
Indexing, leaf-datasets, text width=6em [GNN-ret [44], PG-RAG [45], KGP [46], GraphCoder [47], AVIS[48], SayPlan [49] HotpotQA [50], GraphRetriever [51],
KAR [26], GRAG [52], KET-RAG [53], PIKE-RAG [54], OG-RAG [55], ArchRAG [56], RAPTOR [18], HippoRAG2 [57], text width=50.7em,modelnode-datasets] ] [ Graphs as
@@
Construction from Corpus, leaf-datasets, text width=11em [StructRAG[13], GraphRAG[24], FastRAG [58], QUEST[59], Structure-guided Prompt[60]
LightRAG[25], MedRAG [61], GraphReader [62], LuminiRAG [63], HippoRAG [64], modelnode-datasets, text width=38em] ] [ GraphRAG with Existing
@@
SubGraphRAG [69], PoG [70], LEGO-GraphRAG [71], KnowNet [72], modelnode-datasets, text width=38em] ] ] [ Hybrid
GraphRAG, leaf-datasets, text width=6em [GoR [73], MedGraphRAG [74], KG2RAG [75], HYBGRAG [76], CodexGraph [77], modelnode-datasets, text width=50.7em] ] ] [ Knowledge
@@
CancerKG [79], G-Retriever [80], modelnode-metrics, text width=38em] ] [ Logical-based Retriever, leaf-metrics, text width=11em [RoG [65], RD-P [81], RuleRAG [82], KGL [83], RiTeK [84], modelnode-metrics, text width=38em] ] [ GNN-based Retriever, leaf-metrics, text width=11em [GNN-Ret [44], SURGE [85], GNN-RAG [86], modelnode-metrics, text width=38em] ] [ LLM-based Retriever, leaf-metrics, text width=11em [ToG [21], LightRAG [25], KGP [46], MEG [87], TQA-KG [88], E2GraphRAG [89], modelnode-metrics, text width=38em] ] [ RL-based Retriever, leaf-metrics, text width=11em [KnowGPT [5], Spider [90], GraphRAG-R1 [91], modelnode-metrics, text width=38em] ] ] [ Retrieval
Strategy, leaf-metrics, text width=6em [ Multi-round Retrieval, leaf-metrics, text width=11em [GoR [73], DialogGSR [92], Graph-CoT [93], modelnode-metrics, text width=38em] ] [ Post-retrieval, leaf-metrics, text width=11em [KGR [66], Readi [94], GCR [95], CoK [96], modelnode-metrics, text width=38em] ] [ Hybrid Retrieval, leaf-metrics, text width=11em [StructRAG [13], ToG 2.0 [22], HYBGRAG [76], modelnode-metrics, text width=38em] ] ] ] [ Knowledge
@@
Knowledge, leaf-methods, text width=11em [GRAG [52], RHO [102], GNP [103], InstructGLM [104], LLAGA [105], modelnode-methods, text width=38em] ] ] [ In-context
@@
Graph Refinement, leaf-methods, text width=11em [KELP [67], FMEA-KG [110], EtD [111], PoG [70], CogMG [112], modelnode-methods, text width=38em] ] ] ] ]